# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v2

@@ Appendix A Within Model Feature Space Similarity Results
To examine SAE feature space similarities when trained on the same model, we train multiple SAEs on a one-layer Tinystories model.
@@
SAE Training Parameters. We train two SAEs at different seeds for a TinyStories-1L-21M model (Eldan, 2023). Both SAEs are trained at 30k steps with expansion factor of 16, for a 16384 total features, on a TinyStories dataset, and using ReLU as the activation function. We use a batch size of 4096 tokens and train with an Adam optimizer, applying a constant learning rate of $1\times 10^{-5}$. The learning rate follows a warm-up-free schedule with a decay period spanning 20% of the total training steps. L1 sparsity regularization is applied with a coefficient of 5, and its effect gradually increases over the first 5% of training steps.
We obtain SAE activations using 500 samples from the TinyStories dataset with a max sequence length of 128 (for a total of 64000 tokens).
@@
ManyA-to-1B. This experiment allows multiple features from SAE $A$ to map onto a single feature from SAE $B$. Before filtering non-concept features, this comparison achieves an SVCCA score of 0.73, retaining 9930 feature pairs. After filtering for non-concept features (for this study, we only filter for periods and padding tokens), this pair achieves an SVCCA score of 0.94. Randomly paired achieves a score of 0.004 for 10 runs. We found that filtering for 1-1 did not change the score much.
In contrast, for 10 random runs where we randomly choose 7 tokens that are not part of the non-concept tokens, the filtering of these randomly chosen tokens does not change the unfiltered SVCCA score of 0.73.
@@
Appendix B Cross Model Llama Results
We compare Llama 3-8B and Llama 3.1-8B at each model’s Layer 25 residual stream, a later layer in both models. The Llama 3 SAE has a width of 65536 features and uses Gated ReLU (Rajamanoharan et al., 2024a), while the Llama 3.1 SAE has a width of 32768 features and uses JumpReLU (Rajamanoharan et al., 2024b). We obtain activations using 100 samples with a max sequence length of 100, for a total of 10000 tokens. After filtering, this comparison achieves an SVCCA score of 0.3, comparable to the score of later layers in both Gemma and Pythia.
Appendix C Universal Vs True Features
@@
Appendix D Illustrated Steps of Comparison Method
@@
Appendix E Noise of Many-to-1 mappings
@@
Appendix F Noise of Non-Concept Feature Pairings
@@
Appendix G Concept Categories List and Further Analysis
@@
Appendix H Additional Results for Pythia and Gemma

### v3

@@ E.2 Comparing SAEs trained with different seeds
Tinystories Results . To examine SAE feature space similarities when trained on the same model, we train multiple SAEs on a one-layer Tinystories model.
@@
SAE Training Parameters. We train two SAEs at different seeds for a TinyStories-1L-21M model (Eldan, 2023). Both SAEs are trained at 30k steps with expansion factor of 16, for a 16384 total features, on a TinyStories dataset, and using ReLU as the activation function. We use a batch size of 4096 tokens and train with an Adam optimizer, applying a constant learning rate of $1\times 10^{-5}$. The learning rate follows a warm-up-free schedule with a decay period spanning 20% of the total training steps. L1 sparsity regularization is applied with a coefficient of 5, and its effect gradually increases over the first 5% of training steps. We obtain SAE activations using 500 samples from the TinyStories dataset with a max sequence length of 128 (for a total of 64000 tokens).
@@
ManyA-to-1B. This experiment allows multiple features from SAE $A$ to map onto a single feature from SAE $B$. Before filtering non-concept features, this comparison achieves an SVCCA score of 0.73, retaining 9930 feature pairs. After filtering for non-concept features (for this study, we only filter for periods and padding tokens), this pair achieves an SVCCA score of 0.94. Randomly paired achieves a score of 0.004 for 10 runs. We found that filtering for 1-1 did not change the score much. In contrast, for 10 random runs where we randomly choose 7 tokens that are not part of the non-concept tokens, the filtering of these randomly chosen tokens does not change the unfiltered SVCCA score of 0.73.
@@
E.3 Comparing SAEs with different widths
Figure 10 shows SAEs with 32k latents vs SAEs with 16k latents, both trained on Pythia-70m. This 32k SAE is the SAE shown on the columns (x-axis) of Figure 9. After filtering, around 3.47% of feature pairs are kept on average. Figure 11 shows an SAE with 64k latents vs an SAE with 32k latents, both trained on Pythia-70m. After filtering, around 3.88% of feature pairs are kept on average. The reason why this may occur is discussed in Section §B.
In general, we do not find noticeable trends in scores as we vary SAE widths. This may be because each SAE varies in the features they capture Leask et al. (2025), so while they capture feature similarities across models, the similar features that they do capture are not the same for each SAE. We did not evaluate for feature splitting or absorption (Chanin et al., 2024).
Figure 10: (a) SVCCA 1-1 paired scores of an SAE with 32k latents vs an SAE with 16k latents, (b) RSA 1-1 paired scores of the SAEs.
Figure 11: (a) SVCCA 1-1 paired scores of an SAE with 64k latents vs an SAE with 32k latents, (b) RSA 1-1 paired scores of the SAEs.
E.4 Comparing SAEs trained on different datasets
We compare SAEs on Pythia-70m trained on different training data. We trained an SAE trained on the residual stream layers of Pythia-70m using 100 million tokens using the RedPajama dataset at a context length of 256, with 32768 feature neurons (an expansion factor of 64 and Top-K) of 32. We compare our SAE with Eleuther’s SAE trained at residual streams using 8.2 billion tokens from the Pile training set at a context length of 2049 with 32768 feature neurons and Top-K of 16.
To obtain activations, we use 200 samples with a max sequence length of 200, and after filtering around 4.58% of feature pairs are kept on average, in contrast to the 20 to 30% of feature pairs kept by comparing Eleuther’s pretrained SAEs with each other, perhaps due to the lower quality of our SAEs as discussed in Section §B.
Figure 12 shows that the two sets of SAEs show some similarities, such as at Layers 1, 3, and 5, but also notable dissimilarities, such as having nearly no similarities at Layers 2 and 4. These discrepencies may be due to the difference in training data, as it has been shown that SAEs are highly dataset dependent (Kissane et al., 2024b). However, despite these discrepencies in individual features, the feature subspaces we compare appears to be highly similar, which may suggest that while the specific values captured by feature weight vectors may differ, these values lie in similar latent space regions that capture similar semantics.
Figure 12: (a) SVCCA 1-1 paired scores of Eleuther’s SAE (trained on 8.2 billion tokens from the Pile) vs our SAE (trained on 100 million tokens from RedPajamas). Both have 32k latents. (b) RSA 1-1 paired scores of the SAEs.
We also train SAEs using 100 million tokens from OpenWebText and compare them to the SAEs we trained on RedPajamas. These results are shown in Figure 13. Similar to the results in Figure 12, we find both similarities and dissimlarities which suggest that SAEs are highly dataset dependent.
Figure 13: (a) SVCCA 1-1 paired scores of our SAEs trained on RedPajamas (columns, x-axis) vs our SAEs trained on OpenWebText (rows, y-axis). Both have 32k latents. (b) RSA 1-1 paired scores of the SAEs.
Appendix F Cross Model SAE Width Variation Results
We compare what happens to similarity scores as we increase SAE widths for both Pythia-70m and Pythia-160m. All SAEs were trained using 100 million tokens from RedPajamas and use Top-K of 32. Similar to Section §E.3, we do not find noticeable trends in scores as we vary SAE widths. This may be because each SAE varies in the features they capture, so while they capture feature similarities across models, the similar features that they do capture are not the same for each SAE.
Figure 14 SVCCA and RSA 1-1 paired scores of 24k latent SAEs trained on Pythia-160m vs 16k latent SAEs trained on Pythia-70m, and Figure 15 SVCCA and RSA 1-1 paired scores of 49k latent SAEs trained on Pythia-160m vs 32k latent SAEs trained on Pythia-70m. In both figures, while SVCCA scores are moderately high for some layers, the RSA scores are very low.
Figure 14: (a) SVCCA 1-1 paired scores of 24k latent SAEs trained on Pythia-160m vs 16k latent SAEs trained on Pythia-70m. (b) RSA 1-1 paired scores of the SAEs.
Figure 15: (a) SVCCA 1-1 paired scores of 49k latent SAEs trained on Pythia-160m vs 32k latent SAEs trained on Pythia-70m. (b) RSA 1-1 paired scores of the SAEs.
Appendix G Varying factors used to obtain activations
G.1 Comparing different datasets to obtain activations
We compare results using different datasets, namely the RedPajamas and OpenWebText datasets, to obtain dataset activations used for matching features by activation correlation. For Eleuther’s 32k SAEs trained on Pythia-160m vs on Pythia-70m, Figure 16 displays results from using the RedPajamas dataset, whereas Figure 17 displays results from using the OpenWebText dataset. Overall, the results are similar, with scores from OpenWebText being slightly higher in most cases. This demonstrates that our method is not highly dataset dependent when obtaining activations to pair features by correlation.
Note that these results differ from previous results in Figure 2 because for features matching using Many-to-1 (before filtering), we are using Pythia 160m as the model whose features can have "Many" mappings and Pythia 70m as model whose features can only have "1" mapping, whereas the results in Figure 2 reverse these roles. As implied in Section §J, the results of using a model in the "Many" or "1" places of Many-to-1 are not symmetric, partly due to using max correlation to choose a feature’s partner.
Figure 16: (a) SVCCA 1-1 paired scores of Eleuther’s 32k SAEs trained on Pythia-160m vs on Pythia-70m. (b) RSA 1-1 paired scores of the SAEs. The dataset activations were obtained using 40k tokens from the RedPajamas dataset.
Figure 17: (a) SVCCA 1-1 paired scores of Eleuther’s 32k SAEs trained on Pythia-160m vs on Pythia-70m. (b) RSA 1-1 paired scores of the SAEs. The dataset activations were obtained using 40k tokens from the OpenWebText dataset.
G.2 Comparing more data to obtain activations
Figure 18 shows SVCCA Many-1 paired scores of SAEs for layers in Pythia-70m vs layers in Pythia-160m. This is run with 90k tokens, using 300 samples with a max sequence length of 300. Layer 3 appears to be an outlier, perhaps due to errors in feature matching. These results are similar to and consistent with the results of Figure 20, suggesting that other method is not heavily dependent on the number of tokens used to obtain activations.
Figure 18: (a) SVCCA Many-1 paired scores of SAEs for layers in Pythia-70m vs layers in Pythia-160m. This is run with 90k tokens, using 300 samples with a max sequence length of 300. Layer 3 appears to be an outlier, perhaps due to errors in feature matching. This is consistent with Figure 20.
Appendix H Universal Vs True Features
@@
Stated another way, this paper is not claiming that different SAEs consistently learn the same universal features captured by LLMs. Rather, it is claiming that LLMs learn weakly universal features, and that SAEs can capture not only these features, but their relations. However, one notable limitation is that SAEs cannot capture these features consistently.
Appendix I Illustrated Steps of Comparison Method
@@
Appendix J Noise of Many-to-1 mappings
@@
Appendix K Noise of Non-Concept Feature Pairings
@@
Appendix L Concept Categories List and Further Analysis
@@
Appendix M Additional Results for Pythia and Gemma